2 (b) Prove that + 3 cos(atx) O has at least two solutions with x € (-1,1]. [20 Marks] 1 + x2 (c) State the Rolles Theorem.
Answer #1

☺ Consider the function F: [1] IR de fired by Fras 2 + 3 Costa Ito Now Fro) = 2+3 = 5 > f(1) = 2 + 3cost 2 ے I-3 =-2 lo Hence1-3 - 2 < 0 This feod. f(-1) 2o. Hence One to root dies between otot -1 and 0 on Rolles Theescary: Ler fo[a, ne] – Ik be a fuThen by Rolks therum there exist CE(x, p) Such that f(o)zo- New fm a 42 - 3a sinna (ltage Now Play <ot tf(0,1). Fóra) 7o. X